# DeepMIDE: A Multivariate Spatio-Temporal Method for Ultra-Scale Offshore Wind Energy Forecasting

### v1

@@ 3.1 Local observations from the NY/NJ Bight
Wind speeds are measured using three floating lidar buoys, namely E05 Hudson North (E05N), E06 Hudson South (E06) (New York State Energy Development Authority, 2019), and ASOW6 (Atlantic Shores Offshore Wind, 2021). Figure 1(a) shows the locations of the three sites on top of the designated offshore wind energy areas in the NY/NJ Bight. E05N is the farthest location from shore, and is $\sim$77777777 kilometers (km) away from E06. ASOW6 is the southernmost and closest location to shore. E06 and ASOW6 are $\sim$54545454 km apart. The observations are recorded in $10$-min resolution, for a total duration of approximately $8.5$ months spanning the years of 2020 and 2021, at three heights: $100$m, $140$m, and $180$m. The data from ASOW6 had some missing data streaks, which were imputed using another buoy in close proximity. Figure 1(b) shows the wind rose plot of the wind speed data (pooled across all heights and locations), which shows dominant southwesterly winds.
Figure 1(c) shows the multivariate time series of the wind speeds at all locations and heights during a select period in March 2020, clearly demonstrating strong spatial, temporal, and vertical dependencies. Pairwise Pearson correlation coefficients across the three spatial sites (averaged over height) ranged between $0.853$ (between E05 and ASOW6), to $0.928$ (between E06 and ASOW6), suggesting (perhaps unsurprisingly) strong and distant-dependent spatial correlations. Correlations of wind speeds across multiple heights (averaged over space) are estimated at $0.974$ (between $100$m and $180$m), $0.988$ (between $100$m and $140$m), and $0.992$ (between $140$m and $180$m), suggesting strong height dependencies as functions of the vertical separation. Interestingly, height correlations (at the same vertical separation) appear to be stronger at larger altitudes. Height correlations are more pronounced than those over space, further attesting to the need for modeling the multi-height time series jointly, in contrast to the vast majority of existing space-time wind forecasting methods.
@@ 3.2 Exogenous information from a regional NWP model
A daily real-time version of the Weather Research and Forecasting (WRF) model, called RU-WRF, generates hourly outputs, at $3$-km resolution, for several meteorological variables, including multi-height wind speeds, sea-surface temperature, humidity, and pressure (Dicopoulos et al., 2021). The model is tailored to the U.S. Mid- and North- Atlantic offshore wind energy areas. It has been independently validated by the National Renewable Energy Laboratory (NREL) (Optis et al., 2020) and has been continuously evaluated and improved since then (Dicopoulos et al., 2021). Its spatial domain spans the coastal region from southern Massachusetts to North Carolina. Figure 2 shows RU-WRF outputs, in the form of spatial weather maps, on a select day and time in June, 2021 on top of the offshore wind energy areas in the NY/NJ Bight.
Figure 2: Spatial weather maps from RU-WRF (wind speed at 100/140/180m, sea-surface temperature, humidity, and pressure) on June 11th, 2021 at 8:00 GMT on top of the offshore wind energy areas (denoted by dashed polygons) in the NY/NJ Bight. Red stars denote the spatial locations where local measurements are available.
@@ 3.3 Asymmetry analysis
Due to advection effects, wind fields are known to exhibit strong signs of asymmetry in spatio-temporal dependencies. This means that along-wind dependence (i.e., the dependence between the wind conditions at an upstream location at time $t$ and that at a downstream location at time $t+u$) is expected to be stronger than opposite-wind dependence (i.e., the dependence between the wind conditions at a downstream location at time $t$ and that at an upstream location at time $t+u$) (Gneiting, 2002; Stein, 2005; Ezzat et al., 2018). Thus, adequate models should be able to capture those complex interaction effects. To demonstrate the existence of asymmetry in our data, we first de-trend the time series data at each height and location by fitting a time- and height-specific trend comprising a series of diurnal harmonics, as expressed in (1).
$\small D_{t}^{g}(\mathbf{s})=d_{0}^{g}(\mathbf{s})+d_{1}^{g}(\mathbf{s})\sin\left(\frac{2\pi t}{24}\right)+d_{2}^{g}(\mathbf{s})\cos\left(\frac{2\pi t}{24}\right)+d_{3}^{g}(\mathbf{s})\sin\left(\frac{4\pi t}{24}\right)+d_{4}^{g}(\mathbf{s})\cos\left(\frac{4\pi t}{24}\right),$ (1)
@@
where $D_{t}^{g}(\mathbf{s})$ is the diurnal trend at time $t\in\mathbb{Z}^{+}$, location $\mathbf{s}\in\mathbb{R}^{2}$, and height $g\in\mathbb{Z}^{+}$, whereas $d_{0}^{g}(\mathbf{s}),\ldots,d_{4}^{g}(\mathbf{s})$ are the correspondent regression coefficients. The de-trended data is then used to compute an estimate of asymmetry at the $g$th height, denoted as $a^{g}(\mathbf{s}_{i},\mathbf{s}_{j},u)$, and defined as the difference in spatio-temporal semi-variograms (as a proxy for differences in spatio-temporal covariances). This is expressed in (2).
$a^{g}(\mathbf{s}_{i},\mathbf{s}_{j},u):=\delta^{g}(\mathbf{s}_{i},\mathbf{s}_{j},u)-\delta^{g}(\mathbf{s}_{j},\mathbf{s}_{i},u),$ (2)
@@
where $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ denote the coordinates of the $i$th and $j$th locations, $u$ is the time lag, and $\delta^{g}(\cdot,\cdot,\cdot)$ is the spatio-temporal semi-variogram at height $g$, which is defined as in (3).
$\delta^{g}\left(\mathbf{s}_{i},\mathbf{s}_{j},u\right)=\frac{1}{2(N-u-1)}\sum_{t=1}^{N-u-1}\left\{r^{g}_{t+u}\left(\mathbf{s}_{i}\right)-r^{g}_{t}\left(\mathbf{s}_{j}\right)\right\}^{2},$ (3)
@@
where $N$ is the total number of data points, and $r^{g}_{t}\left(\mathbf{s}_{j}\right)$ denote the value of the de-trended time series at location $\mathbf{s}_{j}$, time $t$, and height $g$.
@@
Figure 3 shows the average values of $a^{g}(\cdot,\cdot,\cdot)$ versus the time lag $u$ (x-axis), for different heights (different columns), partitioned into two wind speed regimes (different rows), representing weak and strong winds respectively (as defined by a threshold of $8$ m/s). A few key observations can be deduced. First, it is clear that there is noticeable levels of spatial-temporal asymmetry in our data, as evident by the departure from the full-symmetry baseline represented by the horizontal dashed line at 0. This asymmetry is the result of the wind propagation across the prevailing southwesterly wind (recall Figure 1(b)). The asymmetry appears to be maximal at the $\sim$1111-$3$-hour time lag, which aligns with the expected time for wind conditions to propagate across the spatial locations considered. Interestingly, asymmetry appears to be stronger with larger altitudes and stronger winds. This means that changes in wind speed over time and height would be accompanied by changes in the strength of asymmetry. An adequate model would not only need to account for the existence of asymmetry, but also for the time- and height-varying dependence structure driven by wind advection dynamics. Existing asymmetric space-time models typically assume a constant level of asymmetry across height and time (Stein, 2005; Gneiting et al., 2006a), or at best, may model some time-varying advection effects, but overlook the vertical variations in asymmetry levels (Ye et al., 2024).
@@ 4 Metholodogy
To model the complex dependencies in the local wind fields across space, time, and height, we propose a multi-output deep IDE model with a multivariate, non-stationary, and state-dependent kernel, for which key physically meaningful parameters are learned using an embedded deep learning model. Our approach is dubbed as the Deep Multivariate Integro-Difference Equation model, or in short, DeepMIDE. We start by a brief background to the IDE model, then present the essence and technical details of DeepMIDE.
@@ 4.1 Background to the IDE framework
Stochastic IDE models are a class of hierarchical space-time models comprised of a data model and a process model (Cressie and Wikle, 2011). The data model relates a latent state process to a set of observable measurements. In the linear-Gaussian case, it can be written as in (4).
@@
where $\mathbf{z}_{t}$ is an $m$-dimensional vector of spatial observations at time $t$, while $\mathbf{y}_{t}$ is the correspondent $n$-dimensional vector of the latent state process at time $t$. The matrix $\mathbf{h}_{t}$ is an $m\times n$ observation mapping matrix, whereas $\boldsymbol{\xi}_{t}$ is an $m$-dimensional vector of measurement errors, which are independent of $\mathbf{y}_{t}$. Assuming locally linear dynamics, the process model describes the conditional first-order evolution of the state process through a spatial integral operator, as expressed in (5).
[…]
@@
where ${y}_{t}(\cdot)$ denotes the spatial process on domain $\mathcal{D}_{s}$ at time $t$. The error process $\nu_{t}(\cdot)$ is independent in time but may be correlated in space, whereas $k(\mathbf{s},\mathbf{x};\boldsymbol{\theta})$ is a spatial redistribution kernel characterized by a set of parameters denoted by $\boldsymbol{\theta}$.
@@
Stochastic IDEs have received significant attention in the space-time modeling literature due to their versatility and ability to integrate mechanistic descriptions of physical and domain knowledge. To date, the overwhelming majority of IDE models have been univariate, i.e. they model a single stochastic process over space and time (Wikle, 2002; Xu et al., 2005; Ke Xu and Fox, 2005; Wikle and Holan, 2011; Calder et al., 2011; Liu et al., 2022; Ye and Ezzat, 2024). However, many real-world processes do not exist in isolation; they are influenced by, and interact with, other concomitant processes. Very little attention has been dedicated to multivariate IDE models, and even less so for wind forecasting applications. In Section 4.2, we propose a multivariate IDE to model and forecast the wind speed over space, time, and height, as motivated by large-scale offshore wind turbines.
@@
4.2 A multivariate IDE model for multi-height space-time offshore wind forecasting
@@
Let $Z^{(g)}_{t}(\mathbf{s})$ be the wind speed observation at location $\mathbf{s}$, time $t$, and height $g$. We denote by $\mathbf{Z}^{(g)}_{t}\equiv(Z^{(g)}_{t}(\mathbf{s}_{1}),\ldots,Z^{(g)}_{t}(\mathbf{s}_{m}))^{\top}$ the $m$-dimensional spatial vector of observations at a particular time and height, whereas $\mathbf{Z}_{t}\equiv(\mathbf{Z}^{(1)}_{t},...,\mathbf{Z}^{(G)}_{t})^{\top}$ is the $mG\times 1$ matrix comprising the observations at time $t$, across all locations and heights, such that $G$ is the number of heights where measurements are available. We express the data model for the multivariate IDE as in (6).
@@
where $\mathbf{Y}_{t}\equiv(\mathbf{Y}^{(1)}_{t},\ldots,\mathbf{Y}^{(P)}_{t})^{\top}$ is the $nP\times 1$ matrix representing the underlying state process at time $t$, across all locations and heights. The matrix $\mathbf{H}_{t}$ is a $mG\times nP$ observation mapping matrix, whereas $\boldsymbol{\epsilon}_{t}$ denotes the zero-mean measurement error vector, which we assume to be a multivariate Gaussian that is independent in time but not in space, with a spatial covariance matrix $\Sigma_{\boldsymbol{\epsilon}}$, for which the entries are determined using a squared exponential covariance function $c_{\epsilon}(d)=\sigma_{\epsilon}^{2}\exp{(-\frac{1}{2}(\frac{\|d\|^{2}}{\ell^{2}_{\epsilon}})})$, where $d$ is the pairwise spatial distance, while $\sigma_{\epsilon}$ and $\ell_{\epsilon}$ are the variance and length-scale parameters, respectively. The Gaussian assumption for the measurement error is non-problematic after an appropriate Box-Cox transformation which makes the wind speeds nearly Gaussian. For our data, the Box-Cox power parameter is estimated at $0.504$, which is almost equivalent to the square root transformations common for modeling wind speeds in the literature (Haslett and Raftery, 1989; Stein, 2005; Gneiting et al., 2006b).
@@
Considering multivariate spatio-temporal processes representing the offshore wind speeds in a space-height continuum, and discrete time, $\{Y_{t}^{(p)}(\cdot)\}$, then the multivariate version of the stochastic IDE process model can be written as in (7).
[…]
@@
such that $k^{(pq)}(\mathbf{s},\mathbf{x};t,\boldsymbol{\theta}_{t}^{(pq)})$ is the redistribution kernel of the multivariate IDE, which determines the contribution of height $q$ at location $\mathbf{x}$ and time $t-1$ in describing the height $p$ at location $\mathbf{s}$ at the current time $t$. The redistribution kernel is assumed to depend on a set of time-varying parameters, denoted by $\boldsymbol{\theta}_{t}^{(pq)}$. The state error vector at time $t$, across all heights and locations, is denoted by $\boldsymbol{\eta}_{t}=(\boldsymbol{\eta}_{t}(\mathbf{s}_{1}),...,\boldsymbol{\eta}_{t}(\mathbf{s}_{n}))^{\top}$ such that $\boldsymbol{\eta}_{t}(\mathbf{s})\equiv(\eta_{t}^{(1)}(\mathbf{s}),...,\eta_{t}^{(P)}(\mathbf{s}))^{\top}$ is the corresponding error vector at location $\mathbf{s}$ and time $t$, across all $P$ heights. The process error term is also assumed to be a zero-mean Gaussian which is independent in time but not in space, with a spatial covariance matrix $\Sigma_{\boldsymbol{\eta}}$, for which the entries are determined using a squared exponential covariance function $c_{\eta}(d)=\sigma_{\eta}^{2}\exp{(-\frac{1}{2}(\frac{\|d\|^{2}}{\ell^{2}_{\eta}})})$, where $\sigma_{\eta}$ and $\ell_{\eta}$ are the variance and length-scale parameters, respectively.
@@
Modeling the redistribution kernel $k^{(pq)}(\mathbf{s},\mathbf{x};t,\boldsymbol{\theta}_{t}^{(pq)})$ is perhaps the most pivotal aspect of IDE models. The asymmetry analysis in Section 3.3 demonstrated the need for capturing time- and height-varying advection dynamics. Here, we follow a similar pursuit to the recent work of Salvaña et al. (2023) who introduced asymmetric cross-covariance functions with multiple advections. We leverage their key theorem to propose a multivariate redistribution kernel, as expressed in (8), that is tailored to model the dependence structure in offshore wind fields, considering time- and height-varying advections.
@@
$k^{(pq)}\left(\mathbf{s},\mathbf{x};t,\boldsymbol{\theta}_{t}^{(pq)}\right)=\exp\left(-\frac{\|\mathbf{s}-\mathbf{x}-\boldsymbol{\theta}_{t}^{p}t+\boldsymbol{\theta}_{t}^{q}(t+1)\|^{2}}{\ell_{\mathbbm{1}_{p=q}}^{2}}\right),$ (8)
@@
where $\ell_{\mathbbm{1}_{(p=q)}}$ $\forall$ $p,q\in\{1,\ldots,P\}$ denote a set of height-dependent diffusion parameters, $\mathbbm{1}_{(\cdot)}$ is the indicator function, and the spatial location $\mathbf{s}$, $\mathbf{x}$ in this paper is vectors containing latitude and longitude. The kernel $k^{(pq)}\left(\mathbf{s},\mathbf{x};t,\boldsymbol{\theta}_{t}^{(pq)}\right)$ is characterized by a set of time- and height-varying advection vectors $\boldsymbol{\theta}_{t}^{(pq)}=\{\boldsymbol{\theta}_{t}^{p},\boldsymbol{\theta}_{t}^{q}\}$, which collectively encode the physics of local wind field formation and propagation, and are assumed to be random in $\mathbb{R}^{2}$.

### v2

@@ 3.1 Local observations from the NY/NJ Bight
Wind speeds are measured using three floating lidar buoys, namely E05 Hudson North (E05N), E06 Hudson South (E06) (measurement), and ASOW6 (asow6). Figure 1(a) shows the locations of the three sites on top of the designated offshore wind energy areas in the NY/NJ Bight. E05N is the farthest location from shore, and is $\sim$$77$ kilometers (km) away from E06. ASOW6 is the southernmost and closest location to shore. E06 and ASOW6 are $\sim$$54$ km apart. The observations are recorded in $10$-min resolution, for a total duration of approximately $8.5$ months spanning the years of 2020 and 2021, at three heights: 100100m, 140140m, and 180180m. The data from ASOW6 had extended periods of missing observations, which were imputed using another buoy in close proximity. Figure 1(b) shows the wind rose plot of the data (pooled across all heights and locations), suggesting dominant southwesterly winds.
Figure 1(c) shows the empirical temporal covariograms (commonly referred to as autocovariance functions) for the offshore wind speeds across all locations and heights. Some clear temporal correlation patterns are observed: (i) For all nine time series, the empirical covariograms show strong correlations at short time lags that decrease with lag but are still distinctly positive at the 24-hour lag; (ii) For different locations (E05N, E06, ASOW6) at the same height, the temporal correlation structure appears to be similar, but with slightly lower overall variability at ASOW6, the site nearest to the coast; (iii) Comparing different heights (100m, 140m, 180m), the covariogram indicates consistently higher variability at greater altitudes (180m) compared to lower altitudes (140m and 100m). Figure 1(d) shows the multivariate time series of the wind speeds at all locations and heights during a select period in March 2020, clearly demonstrating strong spatial, temporal, and vertical dependencies, further motivating the need for a joint modeling approach across the space-time-height trio, as the one advocated in this paper.
@@ 3.2 Exogenous information from a regional NWP model
A daily real-time version of the Weather Research and Forecasting (WRF) model, called RU-WRF, generates hourly outputs, at $3$-km resolution, for several meteorological variables, including multi-height wind speeds, sea-surface temperature, humidity, and pressure (dicopoulos2021weather). The model is tailored to the U.S. Mid- and North- Atlantic offshore wind energy areas. It has been independently validated by the National Renewable Energy Laboratory (NREL) (optis2020validation) and has been continuously evaluated and improved since then (dicopoulos2021weather). Its spatial domain spans the coastal region from southern Massachusetts to North Carolina. Figure 2 shows RU-WRF outputs, in the form of spatial weather maps, on a select day and time in June, 2021 on top of the offshore wind energy areas in the NY/NJ Bight.
Figure 2: Spatial weather maps from RU-WRF (wind speed at 100/140/180m, sea-surface temperature, humidity, and pressure) on June 11th, 2021 at 8:$0$ GMT on top of the planned offshore wind energy areas (denoted by dashed polygons) in the NY/NJ Bight. Red stars denote the spatial locations where local measurements are available.
@@ 3.3 Asymmetry analysis
Due to advection effects, wind fields can exhibit strong signs of asymmetry in spatio-temporal dependencies. This means that along-wind dependence (i.e., the dependence between the wind conditions at an upstream location at time $t$ and that at a downstream location at time $t+u$, where $u>0$ is a time lag) is expected to be stronger than opposite-wind dependence (i.e., the dependence between the wind conditions at a downstream location at time $t$ and that at an upstream location at time $t+u$) (GneitingCov; SteinCov; AzizIEEE). Thus, adequate models should be able to capture those complex interaction effects. To demonstrate the existence of asymmetry in our data, we first de-trend the time series data at each height and location by fitting a time- and height-specific trend comprising a series of diurnal harmonics, as expressed in (1).
$\small D_{t}^{(g)}(\mathbf{s})=d_{0}^{(g)}(\mathbf{s})+d_{1}^{(g)}(\mathbf{s})\sin\left(\frac{2\pi t}{24}\right)+d_{2}^{(g)}(\mathbf{s})\cos\left(\frac{2\pi t}{24}\right)+d_{3}^{(g)}(\mathbf{s})\sin\left(\frac{4\pi t}{24}\right)+d_{4}^{(g)}(\mathbf{s})\cos\left(\frac{4\pi t}{24}\right),$ (1)
@@
where $D_{t}^{(g)}(\mathbf{s})$ is the diurnal trend at time $t\in\mathbb{Z}^{+}$, spatial location $\mathbf{s}\in\mathbb{R}^{2}$ (longitude, latitude), and height $g\in\mathbb{R}^{+}$, whereas $d_{0}^{(g)}(\mathbf{s}),\ldots,d_{4}^{(g)}(\mathbf{s})$ are the corresponding regression coefficients, which are location- and height-dependent. The de-trended data is then used to compute an estimate of asymmetry at the $g$th height, denoted as $a^{(g)}(\mathbf{s}_{i},\mathbf{s}_{j},u)$, and defined as the difference in spatio-temporal semi-variograms (as a proxy for differences in spatio-temporal covariances). This is expressed in (2).
$a^{(g)}(\mathbf{s}_{i},\mathbf{s}_{j},u):=\delta^{(g)}(\mathbf{s}_{i},\mathbf{s}_{j},u)-\delta^{(g)}(\mathbf{s}_{j},\mathbf{s}_{i},u),$ (2)
@@
where $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ denote the coordinates of the $i$th and $j$th locations, and $\delta^{(g)}(\cdot,\cdot,\cdot)$ is the spatio-temporal semi-variogram at height $g$, which is defined as in (3).
$\delta^{(g)}\left(\mathbf{s}_{i},\mathbf{s}_{j},u\right)=\frac{1}{2(N-u-1)}\sum_{t=1}^{N-u-1}\left\{r^{(g)}_{t+u}\left(\mathbf{s}_{i}\right)-r^{(g)}_{t}\left(\mathbf{s}_{j}\right)\right\}^{2},$ (3)
@@
where $N$ is the total number of data points, and $r^{(g)}_{t}\left(\mathbf{s}_{i}\right)$ denotes the value of the de-trended time series at location $\mathbf{s}_{i}$, time $t$, and height $g$, that is, $r^{(g)}_{t}\left(\mathbf{s}_{i}\right)=Z^{(g)}_{t}(\mathbf{s}_{i})-D_{t}^{(g)}(\mathbf{s}_{i})$, such that $Z^{(g)}_{t}(\mathbf{s}_{i})$ is the wind speed observation at location $\mathbf{s}_{i}$, time $t$, and height $g$.
@@
Figure 3 shows the average values of $a^{(g)}(\cdot,\cdot,\cdot)$ versus the time lag $u$ (x-axis), for different heights (different columns), partitioned into two wind speed regimes (different rows), representing weak and strong winds respectively (as defined by a threshold of $8$ m/s). A few key observations can be deduced. First, it is clear that there is noticeable levels of spatial-temporal asymmetry in our data, as evident by the departure from the full-symmetry baseline represented by the horizontal dashed line at 0. This asymmetry is the result of the wind propagation across the prevailing southwesterly wind (recall Figure 1(b)). The asymmetry appears to be maximal at the $\sim$11-33-hour time lag, which aligns with the expected time for wind conditions to propagate across the spatial locations considered. Interestingly, asymmetry appears to be stronger with larger altitudes and stronger winds. This means that changes in wind speed over time and height would be accompanied by changes in the strength of asymmetry. An adequate model would not only need to account for the existence of asymmetry, but also for the time- and height-varying dependence structure driven by wind advection dynamics. Existing asymmetric space-time models typically assume a constant level of asymmetry across height and time (SteinCov; gneiting2006geostatistical), or at best, may model some time-varying advection effects, but overlook the vertical variations in asymmetry levels (ye2023airu; ye2025improved). The above observations warrant a multi-output dynamical forecasting method that can adequately model the complex, time-varying dependence structure found in the multivariate time series data representing the offshore wind speeds across space, time, and height. This will be the focus of Section 4.
@@ 4 Metholodogy
To model the complex dependencies in the local wind fields across space, time, and height, we propose a multi-output deep IDE model with a multivariate, nonstationary kernel, for which key physically meaningful parameters are learned using an embedded deep learning model. Our approach is dubbed as the Deep Multi-output Integro-Difference Equation model, or in short, DeepMIDE. We start by a brief background to the IDE model, then present the essence and technical details of DeepMIDE.
@@ 4.1 Background to the IDE framework
Stochastic IDE models are a class of hierarchical space-time models comprised of a data model and a process model (cressie2011statistics). The data model relates a latent state process to a set of observable measurements. In the linear-Gaussian case, it can be written as in (4).
@@
where $\mathbf{z}_{t}$ is an $m$-dimensional vector of observations at time $t$ and $m$ spatial locations, such that $\mathbf{z}_{t}\equiv(z_{t}(\mathbf{s}_{1}),\ldots,z_{t}(\mathbf{s}_{m}))^{\top}$, while $\mathbf{y}_{t}$ is the correspondent $n$-dimensional vector of the latent state process at time $t$, i.e., $\mathbf{y}_{t}\equiv(y_{t}(\mathbf{s}_{1}^{*}),\ldots,y_{t}(\mathbf{s}_{n}^{*}))^{\top}$. The matrix $\mathbf{h}_{t}$ is an $m\times n$ observation mapping matrix, whereas $\boldsymbol{\xi}_{t}$ is an $m$-dimensional vector of residuals, which are independent of $\mathbf{y}_{t}$. Assuming locally linear dynamics, the process model describes the conditional first-order evolution of the state process through a spatial integral operator, as expressed in (5).
${y}_{t}(\mathbf{s})=\int_{\mathcal{D}_{s}}k(\mathbf{s},\mathbf{x};\boldsymbol{\theta}){y}_{t-1}(\mathbf{x})d\mathbf{x}+\nu_{t}(\mathbf{s}),\quad\mathbf{s}\in\mathcal{D}_{s},$ (5)
@@
where ${y}_{t}(\cdot)$ denotes the spatial process across a region of interest $\mathcal{D}_{s}\in\mathbb{R}^{2}$ at time $t$. The error process $\nu_{t}(\cdot)$ is independent in time but may be correlated in space, whereas $k(\mathbf{s},\mathbf{x};\boldsymbol{\theta})$ is a spatial redistribution kernel characterized by a set of parameters denoted by $\boldsymbol{\theta}$. Discretizing the integral in (5) at $n$ spatial locations yields the following latent process model:
${y}_{t}(\mathbf{s})=\sum_{\mathbf{x}=\mathbf{s}_{1}^{*}}^{\mathbf{s}_{n}^{*}}\tilde{k}(\mathbf{s},\mathbf{x};\boldsymbol{\theta}){y}_{t-1}(\mathbf{x})+\nu_{t}(\mathbf{s}),$} (6)
where $\tilde{k}(\mathbf{s},\mathbf{x};\boldsymbol{\theta})$ is the correspondent redistribution kernel of the discretized process model.
@@
Stochastic IDEs have received significant attention in the space-time modeling literature due to their versatility and ability to integrate mechanistic descriptions of physical and scientific knowledge. To date, the overwhelming majority of IDE models have been single-output, i.e. they model a single stochastic process over space and time (Wikle2002; wikle2005; xu2005; wikle2011; calder2011modeling; liu2022statistical; ye2024ieee). However, many real-world processes do not exist in isolation; they are influenced by, and interact with, other concomitant processes. Very little attention has been dedicated to multi-output IDE models, and even less so for wind forecasting applications. In response, Section 4.2 presents a multi-output IDE to model and forecast the wind energy resource over space, time, and height, as motivated by the emergence of ultra-scale offshore wind turbines.
@@
4.2 A multi-output IDE model for multi-height space-time offshore wind forecasting
@@
For notational consistency, we will use uppercase notation to distinguish between the single-output setting presented in Section 4.1 and the multi-output setting, to be discussed below. Let $Z^{(g)}_{t}(\mathbf{s})$ be the wind speed observation at location $\mathbf{s}$, time $t$, and height $g$. We denote by $\mathbf{Z}^{(g)}_{t}\equiv(Z^{(g)}_{t}(\mathbf{s}_{1}),\ldots,Z^{(g)}_{t}(\mathbf{s}_{m}))^{\top}$ the $m$-dimensional spatial vector of observations at a particular time $t$ and height $g$, whereas $\mathbf{Z}_{t}\equiv(\mathbf{Z}^{(1)}_{t},...,\mathbf{Z}^{(G)}_{t})^{\top}$ is the $mG\times 1$ matrix comprising the observations at time $t$, across the $m$ locations and $G$ heights where measurements are available. We express the data model for the multi-output IDE as in (7).
@@
where $\mathbf{Y}_{t}\equiv(\mathbf{Y}^{(1)}_{t},\ldots,\mathbf{Y}^{(P)}_{t})^{\top}$ is the $nP\times 1$ matrix representing the underlying state process at time $t$, across $n$ locations and $P$ heights. The matrix $\mathbf{H}_{t}$ is a $mG\times nP$ observation mapping matrix, whereas $\boldsymbol{\epsilon}_{t}$ denotes the $mG\times 1$ vector of errors, which are assumed to follow a zero-mean multivariate Gaussian that is independent in time but not necessarily in space, with a spatial covariance matrix $\Sigma_{\boldsymbol{\epsilon}}$, for which the entries are determined using a squared exponential covariance function $c_{\epsilon}(d)=\sigma_{\epsilon}^{2}\exp{(-\frac{1}{2}(\frac{\|d\|^{2}}{\ell^{2}_{\epsilon}})})$, where $d$ is the pairwise spatial distance, while $\sigma_{\epsilon}$ and $\ell_{\epsilon}$ are the variance and length-scale parameters, respectively. The spatial dependence assumption on the error term can be dropped, but is found to slightly benefit the model adequacy by capturing residual spatial dependencies. The Gaussian assumption is non-problematic either after an appropriate Box-Cox transformation applied to the pooled data across all locations and heights, which makes the wind speeds nearly Gaussian. For our data, the Box-Cox parameter is estimated at $0.504$, which is almost equivalent to the square root transformations common for modeling wind speeds in the literature (haslett1989space; SteinCov; gneiting2006calibrated). The forecasts can be retrieved through an inverse Box-Cox transformation directly applied on the model’s outputs. Since the inverse Box-Cox transformation yields the median of the forecast distribution, we followed the approximation described in hyndmanforecasting, which recovers the distribution mean using a second-order Taylor series expansion. The details of the Box-Cox transformation and its inverse are described in the supplementary material document appended to this manuscript.
@@
Considering multi-output spatio-temporal processes representing the offshore wind speeds over space, time, and height, $\{Y_{t}^{(p)}(\cdot)\}$, the multi-output version of the stochastic IDE process model can be written as in (8).
$Y_{t}^{(p)}(\mathbf{s})=\sum_{\mathbf{x}=\mathbf{s}_{1}^{*}}^{\mathbf{s}_{n}^{*}}\sum_{q=1}^{P}k^{(pq)}(\mathbf{s},\mathbf{x};t,\boldsymbol{\theta}_{t}^{(pq)})Y_{t-1}^{(q)}(\mathbf{x})+\eta_{t}^{(p)}(\mathbf{s}),$} (8)
@@
such that $k^{(pq)}(\mathbf{s},\mathbf{x};t,\boldsymbol{\theta}_{t}^{(pq)})$ is the redistribution kernel of the multi-output IDE, which determines the contribution of height $q$ at location $\mathbf{x}$ and time t1t-1 in describing the height $p$ at location $\mathbf{s}$ at the current time $t$. The redistribution kernel is assumed to depend on a set of time-varying parameters, denoted by $\boldsymbol{\theta}_{t}^{(pq)}$. The state error vector at time $t$, across $P$ heights and $n$ locations, is denoted by $\boldsymbol{\eta}_{t}=(\boldsymbol{\eta}_{t}(\mathbf{s}_{1}^{*}),...,\boldsymbol{\eta}_{t}(\mathbf{s}_{n}^{*}))^{\top}$ such that $\boldsymbol{\eta}_{t}(\mathbf{s})\equiv(\eta_{t}^{(1)}(\mathbf{s}),...,\eta_{t}^{(P)}(\mathbf{s}))^{\top}$ is the corresponding error vector at location $\mathbf{s}$ and time $t$, across all $P$ heights. The error term is assumed to be a zero-mean multivariate Gaussian which is independent in time but not necessarily in space, with a spatial covariance matrix $\Sigma_{\boldsymbol{\eta}}$, for which the entries are determined using a squared exponential covariance function $c_{\eta}(d)=\sigma_{\eta}^{2}\exp{(-\frac{1}{2}(\frac{\|d\|^{2}}{\ell^{2}_{\eta}})})$, where $\sigma_{\eta}$ and $\ell_{\eta}$ are the variance and length-scale parameters, respectively.
@@
Modeling the redistribution kernel $k^{(pq)}(\mathbf{s},\mathbf{x};t,\boldsymbol{\theta}_{t}^{(pq)})$ is a critical aspect of IDE models. The asymmetry analysis in Section 3.3 demonstrated the need for capturing time- and height-varying advection dynamics. Here, we follow a similar pursuit to the recent work of madv who introduced asymmetric cross-covariance functions with multiple advections. We leverage their key theorem to propose a multivariate redistribution kernel, as expressed in (9), that is tailored to model the dependence structure in offshore wind fields, considering time- and height-varying advections.
@@
$k^{(pq)}\left(\mathbf{s},\mathbf{x};t,\boldsymbol{\theta}_{t}^{(pq)}\right)={\sigma_{k}^{2}}\exp\left(-\frac{\|\mathbf{s}-\mathbf{x}-\boldsymbol{\theta}_{t}^{p}t+\boldsymbol{\theta}_{t}^{q}(t+1)\|^{2}}{\ell_{\mathbbm{1}_{p=q}}^{2}}\right),$ (9)
@@
where $\sigma_{k}$ is the gain parameter, $\ell_{\mathbbm{1}_{(p=q)}}$ $\forall$ $p,q\in\{1,\ldots,P\}$ denote a set of height-dependent diffusion parameters, $\mathbbm{1}_{(\cdot)}$ is the indicator function, whereas the spatial locations $\mathbf{s}$ and $\mathbf{x}$ are vectors in $\mathbb{R}^{2}$ comprising the latitude and longitude coordinates. The kernel $k^{(pq)}\left(\mathbf{s},\mathbf{x};t,\boldsymbol{\theta}_{t}^{(pq)}\right)$ is characterized by a set of time- and height-varying advection vectors $\boldsymbol{\theta}_{t}^{(pq)}=\{\boldsymbol{\theta}_{t}^{p},\boldsymbol{\theta}_{t}^{q}\}$, which collectively encode the physics of local wind field formation and propagation, and are assumed to be random in $\mathbb{R}^{2}$.